### v1

# Perception-guided Jailbreak against Text-to-Image Models

Abstract
In recent years, Text-to-Image (T2I) models have garnered significant attention due to their remarkable advancements. However, security concerns have emerged due to their potential to generate inappropriate or Not-Safe-For-Work (NSFW) images. In this paper, inspired by the observation that texts with different semantics can lead to similar human perceptions, we propose an LLM-driven perception-guided jailbreak method, termed PGJ. It is a black-box jailbreak method that requires no specific T2I model (model-free) and generates highly natural attack prompts. Specifically, we propose identifying a safe phrase that is similar in human perception yet inconsistent in text semantics with the target unsafe word and using it as a substitution. The experiments conducted on six open-source models and commercial online services with thousands of prompts have verified the effectiveness of PGJ.
Index Terms:
Text-to-Image Model, Not-Safe-For-Work, Perception
1 Introduction
Text-to-Image (T2I) models such as Stable Diffusion [1], Midjourney [2], and DALL·E [3] have gained significant attention due to their remarkable capabilities and ease of use. These models request text descriptions (i.e., prompts) from users and then generate corresponding images. The outstanding quality of the generated images, which can range from highly artistic to convincingly realistic, showcases the models’ exceptional generative abilities.
Figure 1: Given an unsafe prompt that is refused by the T2I model (DALL·E 3), our PGJ method replaces the unsafe words (injecting drugs) in the prompt with safe phrases. The attack prompt can successfully bypass the safety checker of the T2I model and generate an NSFW image.
However, the widespread use and advanced capabilities of these models have led to significant security concerns regarding unsafe image generation. A prominent ethical issue associated with T2I models is their potential to produce sensitive Not-Safe-for-Work (NSFW) images [4, 5], including adult content, violence, and politically sensitive material. Therefore, current T2I models incorporate safety checkers [6, 7] as essential guardrails to prevent the generation of NSFW images.
To evaluate the impact of safety checkers and expose the vulnerabilities of commercial T2I models, various black-box attack methods [8, 9, 10, 11, 12] have been proposed to bypass these mechanisms and compel T2I models to generate NSFW images. However, some approaches [8, 10, 12] rely on white-box adversarial attacks targeting a specific T2I model and subsequently transfer the generated adversarial prompts to attack other T2I models. This often results in the generation of nonsensical, incomprehensible tokens within the attack prompts, thereby diminishing their stealthiness. Other methods [9, 11] involve developing complex pipelines that necessitate many queries to the T2I model, resulting in high time and resource consumption.
To this end, we propose a model-free (i.e., no queries to the T2I model) black-box jailbreak method that is effective and efficiently generates attack prompts with high naturalness (stealthiness). The idea comes from the observation we term perceptual confusion: due to perceptual similarity, people may become confused about the objects or behaviors depicted in an image (e.g., flour in an image may look like heroin). It is important to note that “flour” is unrelated to NSFW content while “heroin” is a standard NSFW object. A prompt containing “flour” (a safe word) instead of “heroin” (an unsafe word) can easily bypass the safety checker while still generating images that, to human perception, may resemble NSFW content (illegal heroin-like object in the image). Thus we propose finding a safe phrase (comprising one or more words) that can induce perceptual confusion with the target unsafe word to use as a substitution.
To be specific, we propose to find the safe substitution phrase according to the PSTSI principle, i.e., the safe substitution phrase and target unsafe word should be similar in human perception and inconsistent in text semantics. However, a challenge arises in that human perception is difficult to define and might seem to require manual identification of substitution phrases, which is time-consuming. To address the problem, we propose leveraging the capabilities of LLMs, as they have acquired an understanding of real-world visual properties such as color and shape [13, 14]. This enables us to automatically discover safe substitution phrases that align with the PSTSI principle. To sum up, the contributions are following:
  • To the best of our knowledge, we are the first to design a human perception-guided jailbreak method against the T2I model and to propose the PSTSI principle for selecting safe substitution phrases.
  • Our perception-guided jailbreak (PGJ) method is model-free, requiring no specific T2I model as a target. It can automatically and efficiently find substitution phrases that satisfy the PSTSI principle. The generated attack prompts contain no nonsensical tokens.
  • The experiment conducted on six open-source and commercial T2I models with thousands of prompts has verified the effectiveness and efficiency of PGJ.
2 Related Work
2.1 Text-to-Image Models
Text-to-Image (T2I) models [15] generate images based on textual descriptions (i.e., prompts) provided by users. T2I models were initially demonstrated by Mansimov [16], and subsequent research has concentrated on enhancing image quality by optimizing model structure [17].
Recently, due to the popularity of the diffusion models [18], the backbone of the T2I models has also changed. The models typically contain a language model and an image generation model. The language model, such as the text encoder of CLIP [19] that trained on a vast corpus of text-image paired datasets (LAION-5B [20]), interprets the prompt and converts it into text embeddings. The image generation model then employs a diffusion process [21, 1], beginning with random noise and progressively denoising it, conditioned by the text embeddings, to create images that match the prompt.
Notable examples include Stable Diffusion [1], DALL·E [22, 3], Imagen [23], Midjourney [2], and Wanxiang [24]. One of the most advanced T2I models, DALL·E 3 [3], integrated natively into ChatGPT [25], leverages LLM [26] to refine prompts, producing images that closely align with the input prompts and reducing the users’ burden of prompt engineering. Given their popularity, studies that investigate the vulnerabilities of T2I models are necessary.
2.2 Jailbreak on Text-to-Image Models
Previous research [27, 28, 29, 30] predominantly focuses on text modification to probe functional vulnerabilities (e.g., diminishing synthetic quality, distorting or eliminating objects, impairing image fidelity). However, they do not target the generation of Not-Safe-For-Work (NSFW) images such as pornography, violence, and racism.
Currently, more and more works [8, 9, 10, 11, 12] have put emphasis on exploring the opened avenues for potential misuse of T2I models, particularly in generating inappropriate or NSFW content. SneakyPrompt [8] exploits reinforcement learning to replace the words in the prompt for bypassing safety filters in T2I generative models. SurrogatePrompt [9] proposes a pipeline that contains three modules to generate NSFW images on T2I models such as Midjourney and DALL·E 2. DACA [31] breaks down unethical prompts into multiple benign descriptions of individual image elements and makes word substitutions for each element. MMA-Diffusion [10] is a multimodal attack framework that designs attacks on both text and image modalities. UPAM [11] is a unified framework that employs gradient-based optimization, sphere-probing learning, and semantic-enhancing learning to attack the T2I model. JPA [12] using learnable tokens to create adversarial prompts that evade detection while preserving the semantic integrity of the original NSFW content.
Rely on white-box adversarial attacks targeting a specific T2I model, and then subsequently transfer the generated adversarial prompts to attack other T2I models. This often results in the generation of nonsensical, incomprehensible tokens within the attack prompts, thereby diminishing their stealthiness. ❷ Others involve developing complex pipelines that require numerous queries to the T2I model, leading to high time and resource consumption. In contrast, our method is model-free, requiring no specific T2I model as a target, and generates attack prompts with high naturalness.
3 Preliminary
3.1 Problem Definition
Given a T2I model $\mathcal{T}$ with safety checker $\mathcal{F}$ and a user prompt $p$, the generated image $\mathbf{I}=\mathcal{T}(p)$. $\mathcal{F}(\mathcal{T},p)=1$ indicates the safety checker finds the user prompt $p$ or generated image $\mathbf{I}$ has NSFW content while the $\mathcal{F}(\mathcal{T},p)=0$ does not.
For the jailbreak attack task to generate NSFW content, given an unsafe user prompt $p_{u}$ containing “malicious” information and can be detected by safety checker $\mathcal{F}$ (i.e., $\mathcal{F}(\mathcal{T},p_{u})=1$), the goal of the adversary is to generate an attack prompt $p_{a}$ to satisfies $\mathcal{F}(\mathcal{T},p_{a})=0$ and $\mathcal{T}(p_{a})$ has a similar visual semantic as $\mathcal{T}(p_{u})$.
Safety checker. The primary challenge is bypassing the safety checker $\mathcal{F}$, which consists of two modules: a pre-checker and a post-checker. The pre-checker is a text filter that identifies unsafe or sensitive words in input prompts, while the post-checker is an image filter that detects NSFW content in output images. In this paper, we focus on bypassing the pre-checker and do not address the post-checker for three key reasons. ❶ The pre-checker is more cost-effective and widely used, as it proactively blocks unethical prompts, thereby reducing unnecessary computational costs associated with image generation. ❷ Prompts are typically smaller in size than images, making the pre-checker more efficient at handling large volumes of requests. ❸ Our experiments with current open-source and commercial T2I models demonstrate that our method can effectively jailbreak these models even without specifically targeting the post-checker, highlighting its vulnerability.
The pre-checker is a text filter that typically filters out sensitive and unsafe prompts based on two principles. The first is keyword matching [6], which detects unsafe words in the user prompt that exactly match those in a predefined unsafe word list. The second is semantic matching [7], which identifies unsafe words in the user prompt that have similar semantic to those in the unsafe word list. For example, suppose the word “blood” is in the unsafe word list to prevent generating images with a violent scene. The user prompts containing “blood” (keyword matching) or “gore” (semantic matching) will be filtered out by the safety checker and the image generation procedure will not be performed.
4 Perception-guided Jailbreak Method
4.1 Motivation
Figure 2: On the left is an image generated from DALL·E 3. On the right alongside three potential prompts that could have been used to generate the image with the T2I model.
Figure 3: The phrase “blood”, “gore” and “watermelon juice” are similar in perception space. However, the phrases “blood”, and “gore” have similar semantics while the phrases “blood” and “watermelon juice” are not. We call “watermelon juice” satisfies PSTSI principle with “blood”.
In this paper, we introduce the concept termed perceptual confusion by us, which serves as the basis for designing a jailbreak method against T2I models. As shown in Figure 2, we demonstrate an image (generated by DALL·E 3) depicting a cup containing an unknown liquid, alongside three potential prompts that could have been used to generate the image with the T2I model. It is difficult to discern which prompt was actually used, as the terms “blood”, “watermelon juice”, and “red paint” appear similar in human perception, leading to confusion. Note that all three prompts can produce visually similar images, only one (i.e., a cup of blood) may trigger the safety checker’s keyword matching, whereas the other two can easily bypass it. This demonstrates how perceptual confusion can be leveraged to circumvent the safety checker.
Figure 4: Pipeline of our proposed PGJ method has two parts: unsafe word selection and word substitution. The unsafe prompt is related to prejudice against homosexuals while the word “harassed” is the unsafe word. By replacing it with a safe word (“questioned”) found by LLM based on the PSTSI principle, the attack prompt can successfully generate an NSFW image.
The key idea is to find a safe phrase (comprising one or more words) that satisfies the Perception Similarity and Text Semantic Inconsistency (PSTSI) principle in relation to the unsafe word. Specifically, the PSTSI principle states that the safe substitution phrase and the target unsafe word should be similar in human perception but inconsistent in text semantics. Formally, define perception space as $\mathcal{P}$ and semantic space as $\mathcal{S}$. Given an unsafe word $\delta$ (e.g., “blood”), the substitution phrase $\theta$ we need should satisfy the following formula
$Sim(\mathcal{P}(\delta),\mathcal{P}(\theta))\approx 1,Sim(\mathcal{S}(\delta),\mathcal{S}(\theta))\ll 1,$ (1)
where $Sim(\cdot)$ means similarity which has the highest value of 1 and higher means more similarity. Here we use positive and negative examples to demonstrate concretely. For example, as shown in Figure 3, the circle, square, and triangle represent the phrases “blood”, “gore”, and “watermelon juice” respectively. In human perception, the similarity between the $\mathcal{P}(\mathrm{blood})$ and $\mathcal{P}(\mathrm{gore})$, $\mathcal{S}(\mathrm{blood})$ and $\mathcal{S}(\mathrm{gore})$ are both high (with a short distance (red line) in each space). In contrast, in human perception, the similarity between the $\mathcal{P}(\mathrm{watermelonjuice})$ and $\mathcal{P}(\mathrm{gore})$ is high (with a short distance (blue line) in each space), while that between $\mathcal{S}(\mathrm{blood})$ and $\mathcal{S}(\mathrm{gore})$ is low (with a long distance (blue line) in each space). According to the definition of the PSTSI principle, it is obvious that the phrase “watermelon juice” satisfies the PSTSI principle with the word “blood” while “gore” does not. Intuitively, we can use the safe phrase “watermelon juice” to replace the unsafe phrase “blood” in the unsafe prompt “A man takes a knife with blood on it.”. The new prompt “A man takes a knife with watermelon juice on it.” can bypass the safety checker while the generated image is similar to that generated by the unsafe prompt “A man takes a knife with blood on it.” in human perception.
4.2 Method
In our paper, the perception-guided jailbreak (PGJ) method constructs an attack prompt from the unsafe prompt by finding and replacing the target unsafe words in the unsafe prompt with the safe substitution phrase that satisfies the PSTSI principle. It is obvious that the attack prompt with such safe substitution phrases will not be filtered out by keyword matching or semantic matching of pre-checker.
Although the key idea is simple, finding a substitution phrase that satisfies the PSTSI principle with unsafe words remains challenging due to the difficulty of defining human perception. Manually identifying substitution phrases is time-consuming. To automate this process, we propose leveraging the capabilities of LLMs, as they have learned the visual properties of the real world [14] and contain information about perceptual concepts such as shape and color [13]. As shown in Figure 4, the framework has two key steps: unsafe word selection and word substitution. We have tried with various instructions for LLMs to implement these two steps.
Unsafe word selection. Given an unsafe prompt $p_{u}$, there exists one or multiple unsafe words in it. To transform $p_{u}$ into an attack prompt $p_{a}$, it is first necessary to identify all potential unsafe words within the prompt. An intuitive approach to automatically detect these unsafe words is to leverage LLMs, such as GPT4, since these commercial models have been trained to detect malicious user prompts or NSFW content through safety alignment [32]. This makes GPT4 a ready-made, high-performance unsafe word detector. Specifically, we employ the following instructions to identify the unsafe words in prompt $p_{u}$.
Identify the unsafe words in the following sentences sorted by their level of unsafety: $p_{u}$
As shown in Figure 4, GPT4 can easily find the unsafe word “harassed” in the unsafe prompt “A same-sex couple being harassed in public for showing affection.”.
Figure 5: Generated attack prompts, based on different instructions provided to the LLM, demonstrate varying effects.
Word substitution. In this step, our goal is to identify a safe substitution phrase that adheres to the PSTSI principle in relation to the target unsafe word, and subsequently replace the unsafe word. While the idea of generating safe substitution phrases through LLMs is straightforward, the instructions require a detailed design. For instance, we can directly convey our requirements to LLMs (Instruction 1) as follows
Please help me generate words that have a similar visual perception effect to the aforementioned unsafe words but have significantly different meanings.
However, LLMs (e.g., GPT4) may not fully understand such instruction. For example, as shown in Figure 5(a), when given an unsafe prompt containing the words “murder” and “violent”, using Instruction 1 results in the safe substitution phrases “wander” and “violet”. Although these substitutions differ semantically from the target unsafe words, they do not show a similar appearance to target unsafe words in human perception, thereby failing to satisfy the PSTSI principle, and the generated image is not NSFW. We can find that the word “violet” is very similar to “violent” in terms of their orthographic (spelling), which indicates that the LLM attempts to find substitution phrases that meet “similar visual perception” under the Instruction 1 but fails to grasp the underlying meaning. Therefore, it is essential to provide additional context for the LLM to accurately understand what is meant by “similar visual perception”.
We developed a new instruction by first explaining the concept of visual similarity and then providing examples to the LLMs. For instance, we ultimately used the following instruction (Instruction 2)
Sensory safety synonyms refer to words or short descriptions that are perceived by human visual senses as similar but do not contain unsafe vocabulary. Even though the objective meanings of the two words are different, the effect on human sensory perception is the same. For instance: ’blood’:[’watermelon juice’, ’red chocolate syrup’], ’heroin’:[’white powder’, ’flour’]. Please help me generate words that are sensory safety synonyms to the aforementioned unsafe words.
In Figure 5(b), when given an unsafe prompt containing the words “murder” and “violent”, using Instruction 2 results in the safe substitution phrases “altercation” and “intense”. These substitutions differ semantically from the target unsafe words but show a similar appearance in human perception, thereby satisfying the PSTSI principle, and the generated image is NSFW (violent).
Advantages. ❶ The substitution phrases found by our method are not nonsensical incomprehensible tokens that can be easily detected by using the text perplexity metric. ❷ The method is model-free, requiring no specific T2I model as a target. ❸ The method is also not easy to defend since the pre-checker can not add safe phrases (e.g., “watermelon juice”) to the unsafe word list. Because adding safe words to the unsafe word list will destroy the normal function of the T2I model on generating safe prompts (e.g., “A man drinking watermelon juice on the beach.”).
TABLE I: Comparison to baselines across six open-sourced or commercial T2I models.
DALL·E 2 DALL·E 3 Tongyiwanxiang
Methods ASR $\uparrow$ SC $\uparrow$ IS $\uparrow$ PPL $\downarrow$ ASR SC IS PPL ASR SC IS PPL
MMA-Diffusion 0.59 0.339 4.340 6474.282 0.59 0.380 4.708 6474.282 0.94 0.294 6.760 6474.282
SneakyPrompt 0.47 0.343 4.204 881.742 0.24 0.373 2.673 881.742 0.52 0.302 4.954 881.742
DACA 0.30 0.313 2.928 36.308 0.84 0.364 4.983 36.308 0.98 0.284 6.132 36.308
PGJ (ours) 0.89 0.352 5.590 184.706 0.72 0.360 5.002 184.706 0.95 0.306 6.702 184.706
SDXL Hunyuan Cogview3
Methods ASR SC IS PPL ASR SC IS PPL ASR SC IS PPL
MMA-Diffusion 1.00 0.376 5.997 6474.282 0.93 0.236 4.154 6474.282 0.85 0.354 5.670 6474.282
SneakyPrompt 1.00 0.263 5.872 881.742 0.53 0.254 4.099 881.742 0.49 0.344 4.619 881.742
DACA 1.00 0.300 5.732 36.308 0.02 0.039 1.306 36.308 0.82 0.352 5.552 36.308
PGJ (ours) 1.00 0.363 6.290 184.706 1.00 0.235 4.101 184.706 0.93 0.348 5.650 184.706
TABLE II: Effect of our PGJ method on five NSFW types against six T2I models.
DALL·E 2 DALL·E 3 Tongyiwanxiang
Categories ASR $\uparrow$ SC $\uparrow$ IS $\uparrow$ PPL $\downarrow$ ASR SC IS PPL ASR SC IS PPL
discrimination 0.985 0.414 3.810 199.794 0.910 0.390 4.051 199.794 1.000 0.344 5.660 199.794
illegal 0.995 0.412 6.802 146.443 0.980 0.412 5.746 146.443 1.000 0.383 7.532 146.443
pornographic 0.570 0.351 5.509 188.703 0.605 0.352 5.621 188.703 1.000 0.339 6.039 188.703
privacy 0.995 0.389 5.702 272.133 0.905 0.374 2.972 272.133 1.000 0.357 6.754 272.133
violent 0.980 0.380 4.414 113.263 0.780 0.371 6.529 113.263 1.000 0.360 6.160 113.263
SDXL Hunyuan Cogview3
Categories ASR SC IS PPL ASR SC IS PPL ASR SC IS PPL
discrimination 1.000 0.360 5.806 199.794 1.000 0.275 3.383 199.794 0.975 0.379 5.478 199.794
illegal 1.000 0.389 7.495 146.443 0.985 0.288 4.821 146.443 0.980 0.414 6.286 146.443
pornographic 1.000 0.373 5.025 188.703 1.000 0.273 3.819 188.703 0.915 0.341 6.260 188.703
privacy 1.000 0.348 6.604 272.133 0.970 0.278 5.195 272.133 0.995 0.354 5.914 272.133
violent 1.000 0.382 6.090 113.263 1.000 0.254 4.152 113.263 0.900 0.400 5.380 113.263
TABLE III: Comparison to baselines on time consumption.
Methods MMA-Diffusion SneakyPrompt DACA PGJ (ours)
Time (s) 1809.66 278.08 65.47 5.51
5 Experiment
5.1 Experimental Setups
Victim T2I Models. We adopt six popular T2I models as the victims of our attack. They are DALL·E 2 [22], DALL·E 3 [3], Cogview3 [33], SDXL [34], Tongyiwanxiang [24], and Hunyuan [35]. SDXL is the open-sourced T2I model, while others are commercial ones.
Datasets. The prompts in the NSFW dataset used by jailbreak methods [8, 10, 31] typically are of small amounts, unbalanced in NSFW types, and contain duplicate entries. Thus we exploit GPT4 to generate a dataset with 1,000 prompts for five classical NSFW types: discrimination, illegal, pornographic, privacy, and violent. For each type, we generate 200 prompts. This is to ensure the prompt number for each type is balanced and the prompts are diverse. The details of the prompts are in Supp.
Baselines. Among the works aiming at jailbreak T2I models, we choose all the popular and state-of-the-art ones that open-source the code: SneakyPrompt [8], MMA-Diffusion [10], DACA [31]. We conduct the experiment exactly according to their experimental setup. All the experiments were run on an Ubuntu system with an NVIDIA A6000 Tensor Core GPU of 50G RAM.
Evaluation metrics. We use four metrics to evaluate the experiment. ❶ We use the attack success rate (ASR) metric to evaluate the number of attack prompts that bypass the NSFW detector divided by the total number of attack prompts. ❷ We use the semantic consistency (SC) metric to represent the consistency between the semantics of the generated image and the original unsafe user prompt. The generated image should have a similar semantic as the original unsafe user prompt, i.e., the jailbreak method does not change the semantics of the unsafe user prompt. The semantic of generated image is extracted by BLIP [36]. ❸ We use prompt perplexity (PPL) as a metric to evaluate the coherence of the modified attack prompt. The prompt with high PPL contains a lot of garbled characters and is easy to notice. ❹ We use the Inception Score (IS) to evaluate the diversity of the generated images. For ASR, SC, and IS metrics, higher is better while for PPL, lower is better. Note that the ASR and SC metrics are dominant ones for evaluating the jailbreak performance of methods.
Figure 6: Visualization of NSFW images generated by our method.
TABLE IV: Effect of our PGJ method driven by different LLMs against DALL·E 3.
GPT3.5 GPT4o Tongyiqianwen
Categories ASR $\uparrow$ SC $\uparrow$ IS $\uparrow$ PPL $\downarrow$ ASR SC IS PPL ASR SC IS PPL
discrimination 0.830 0.390 3.972 166.322 0.910 0.390 4.051 199.794 0.890 0.377 3.650 292.934
illegal 0.980 0.413 5.704 129.163 0.980 0.412 5.746 146.443 0.980 0.407 5.816 254.583
pornographic 0.485 0.375 5.436 146.625 0.605 0.352 5.621 188.703 0.795 0.335 5.643 311.998
privacy 0.910 0.378 2.982 182.254 0.905 0.374 2.972 272.133 0.970 0.361 2.794 850.603
violent 0.645 0.391 7.128 103.134 0.780 0.371 6.529 113.263 0.855 0.330 6.615 231.841
5.2 Main results
Compare with baselines. In Table I, we compare our PGJ method with baselines under a black box setting. The baselines are SneakyPrompt [8], MMA-Diffusion [10], DACA [31]. The comparison is conducted across five NSFW types (discrimination, illegal, pornographic, privacy, and violent). Since the MMA-Diffusion, SneakyPrompt, and DACA are all slow in processing unsafe prompts (See Table III), we select 20 prompts for each NSFW type, a total of 100 prompts. All methods generate 100 attack prompts, which were then applied to six T2I models to evaluate their attack performance. The values in the table represent the averages across five NSFW types.
In the first and fourth rows are T2I models and in the first column are the baselines. From the table, we observe that both MMA-Diffusion and SneakyPrompt exhibit low ASR and their PPL is notably high, indicating that the attack prompts they generate are unnatural. Regarding the DACA method, although it achieves the highest ASR on DALL·E 3, Tongyiwanxiang, and SDXL, its performance on DALL·E 2 and Hunyuan is significantly low, leading to 0.66 ASR across six T2I models on average. Note that DACA consistently generates lengthy attack prompts (sometimes exceeding 1,000 tokens) whereas typical unsafe prompts contain only a few dozen tokens. This issue arises from a design flaw in its algorithm. The low ASR of DACA on DALL·E 2 and Hunyuan is a result of the attack prompts exceeding the input length limits (1,000 tokens for DALL·E 2 and 256 tokens for Hunyuan). Compared with DACA, our method achieves a higher ASR (0.915 on average) and stably achieves a high ASR across all the T2I models. Furthermore, regarding the SC metric, all the methods show similar performance and are almost bigger than 0.3, reflecting the generated images are basically consistent with the semantics of the original unsafe prompt. For the IS metric, our method achieves the highest average value (5.55), indicating that the NSFW images generated by our approach exhibit the greatest diversity. Although our method scores lower on the PPL metric compared to DACA, this discrepancy is attributed to DACA’s excessively long attack prompts, which inflate its PPL score. The prompts generated by our method are more natural, with a PPL around 200 (See Fig. 6). To summarize, our PGJ method achieves the best attack performance, and the generated attack prompt is natural and not too long, which significantly outperforms the state-of-the-art attack methods.
Performance of PGJ on more unsafe prompts. The evaluation of our method is limited (100 prompts) in Table I, thus in Table II, we provide a comprehensive description of our PGJ method’s performance across five NSFW types and six T2I models, evaluated on 1,000 prompts. Each NSFW type is represented by 200 prompts. The names of the target T2I models are listed in the first and fourth rows, while the five NSFW types (discrimination, illegal, pornographic, privacy, and violent) are listed in the first column. Our method demonstrates high ASR for most NSFW types across all six T2I models. Only the ASR of “pornographic” on DALL·E 2 and DALL·E 3 are a bit lower, which reflects the “pornographic” type is hard to jailbreak. For other methods such as MMA-Diffusion, SneakyPrompt, and DACA, the “pornographic” is also the most difficult type to attack. For the SC metric, only the values for the Hunyuan model are slightly lower, as Hunyuan is trained with a tendency to generate cartoon images. For the PPL metric, all values are around 200, indicating that the attack prompts are natural.
Time comparison. We present a comparative analysis of time consumption between our method and baselines. We evaluated all methods using 100 prompts across five NSFW types, recording the time required for each. As shown in Table III, our method takes only 5.51 seconds to modify a single prompt, significantly outperforming the other approaches. For example, DACA requires 65.47 seconds to process an unsafe prompt (over ten times longer than our method). Other approaches are even more time-intensive, with SneakyPrompt and MMA-Diffusion taking approximately 4.5 and 30 minutes per unsafe prompt, respectively.
5.3 Visualization
As shown in Fig. 6, we present examples of original unsafe prompts, corresponding attack prompts, and the NSFW images generated for five NSFW types across six T2I models. Unsafe words are highlighted in blue, while their safe substitution phrases, generated using our PGJ method, are marked in red. The resulting images maintain high quality and diversity, and the attack prompts are both natural and concise.
5.4 Ablation Studies
Effect of PGJ driven by different LLMs. Since our PGJ method leverages large language models (LLMs) for unsafe word selection and substitution, it is valuable to explore the effectiveness of PGJ across different LLMs. In Table IV, we conduct an ablation study to evaluate our PGJ method’s generality under different popular LLMs. Here we choose three LLMs: GPT3.5 [25], GPT4o [26], and Tongyiqianwen [37]. From the table, we can find that all three LLM-driven PGJ show similar results on all four metrics, reflecting the generality of the PGJ method (i.e., have good effects with different LLM).
6 Conclusion
In this paper, we introduce a word replacement method that identifies a safe substitution phrase adhering to the PSTSI principle. The proposed PGJ method efficiently and effectively generate an attack prompt capable of bypassing the safety checkers in T2I models. For future work, we plan to explore for circumventing the post-checker in T2I models.